# From Challenges and Pitfalls to Recommendations and Opportunities: Implementing Federated Learning in Healthcare

### v1

@@ 5.2.1 Unclear in Data Standardization and Harmonization
Healthcare data are often collected and stored in diverse and proprietary formats that do not always adhere to international standards and terminologies, complicating data linkage and reuse. For example, structured clinical data usually contains features that vary with differences in clinical practice across institutions 124, such as diabetes diagnosis, which can involve different glucose measurement methods with varying cut-off points, resulting in hidden heterogeneity that may be overlooked in subsequent statistical analyses. A crucial step before implementing FL in healthcare is to ensure data standardization, harmonization, and interoperability across different cohorts, which are key to the success of FL (Figure 6).
Most simulation studies processed data centrally and generate artificially partitioned datasets without considering the distributed nature of various data silos. This oversight extends to the lack of discussion on how datasets at each client are curated for use in experiments. Despite this, almost all FL frameworks assume the input data is preformatted for model training or preprocessing pipelines. This assumption leads to significant frustration and delays, as the burden of data export and conversion typically falls on clinical data managers who may lack the necessary budget and training. Moreover, among the included studies, only two performed quality or integrity checks on the data. 125 excluded samples with impossible values (e.g., negative heart rates) and inconsistent feature values, while 126 used Principal Component Analysis to filter out noise. Few studies addressed structural or informative missingness, with only 57 and 38 considering imputation methods while also deleting features with high missingness rates. Poor quality imputation and handling of non-random missingness can bias model training.
Figure 6: Data Standardization and Harmonization.
@@
  • Use Consistent Protocols: Collaborating clients must use consistent preprocessing protocols to standardize data to a Common Data Model (CDM) such as the Observational Medical Outcomes Partnership (OMOP) 127. Harmonizing healthcare data to a CDM like OMOP ensures it is interoperable with other clinical datasets, enabling effective merging and analysis across distributed sources and platforms 93.
@@ 5.2.2 Issues in Data Partition
Among all included studies, only a few leveraged natural splits to replicate data collection processes across different hospitals or institutions. For instance, 30 employed the CheXpert dataset, 128 worked with the chest X-ray dataset and 39 extracted metadata from Tissue Source sites in the TCGA dataset for their studies. These datasets naturally reflect the heterogeneity found in real-world clinical data across hospitals and institutions, making them more suitable for FL studies in healthcare 129, 58.
Simulation studies typically used heuristics to artificially create heterogeneous data partitions from a pooled dataset, assigning these partitions to simulated clients, as illustrated in Figure 4. Common synthetic partitioning methods for classification tasks include assigning samples from a limited number of classes to each client, using Dirichlet distribution sampling on class labels, and employing the Pachinko Allocation Method (PAM) when labels have a hierarchical structure 4. For regression tasks, Gaussian Mixture clustering based on t-SNE feature representations has been used to partition datasets among clients 130.
Nonetheless, synthetic partitioning methods may not accurately reflect the intricate heterogeneity found in real-world scenarios 90. Examples from digital histopathology illustrate the limitations of synthetic partitioning methods 131. In digital histopathology, tissue samples are extracted, stained, and digitized, leading to data heterogeneity due to factors such as patient demographics, staining techniques, physical slide storage methods, and digitization processes. Although advancements in staining normalization have reduced some heterogeneity, other sources remain challenging to replicate synthetically, and some may even be unknown 132. These underscore the necessity of conducting cohort experiments with natural splits to ensure that FL models are robust across varied clinical settings. This issue also extends to other areas, including radiology, dermatology, and retinal image analysis.
@@
Another significant issue is the lack of clear definitions and descriptions for train and test set partitions across clients in many studies. 84% included studies failed to explicitly define how these partitions are handled for each client, leading to potential ambiguity in evaluating model performance. This concern is particularly critical in the context of personalized FL, where each client’s test set should be unique to accurately reflect individual data distributions.
@@
  • Complement Simulation Studies with Real-World Data Evaluations: While simulation studies using artificially partitioned datasets can provide valuable insights, it is essential to validate these findings through evaluations on real-world, naturally partitioned datasets. This multi-stage evaluation process ensures that models are tested in both controlled environments and realistic deployment scenarios, improving their generalizability and robustness 133.
@@
  • Incorporate Multiple Types of Skew: Researchers should broaden the scope of their synthetic partitioning strategies to include not just quantity skew but also label skew and feature skew. This would create a more realistic representation of the heterogeneity found in real-world datasets, allowing FL models to be more robust and generalizable across diverse clinical settings 134.
  • Retaining Client-Specific Test Sets: Consider the non-IID nature of data in FL, we suggest that researchers retain a portion of data within each client as a dedicated test set rather than relying on a single, global test set for all clients. This approach provides a more accurate and reliable evaluation, reflecting the unique data distributions of each client, which is particularly important in personalized FL scenarios 133.
@@ 5.3.1 Inadequate Model Selection and Development
  • Standardization for Research and Benchmarking: There should be an emphasis on standardizing model architectures to enable better comparison and benchmarking across studies, especially when trying to establish baseline performance or when deploying FL in more homogeneous environments. Researchers should also consider developing more lightweight models that can be deployed in resource-constrained environments without sacrificing performance.
  • Exploring Simpler, More Interpretable Models: Researchers should not overlook simpler, traditional ML models, particularly in scenarios where interpretability is crucial. These models can be more practical and equally effective in certain healthcare applications, providing a balance between performance and interpretability 135, 21.
  • Enhancing Personalization in FL: More work is needed to develop and refine personalized FL approaches that can adapt to the diverse data distributions encountered in healthcare. This could involve hybrid models that combine the strengths of both global and local models, as well as more sophisticated techniques for model adaptation 49.
  • Addressing Model Heterogeneity: Given the diverse requirements and constraints across different institutions, it is crucial to develop FL strategies that can effectively manage model heterogeneity. This includes exploring federated ensemble learning methods, which allow the aggregation of heterogeneous models and can lead to more robust and accurate predictions 136.
  • Improving Scalability and Efficiency: Future studies should prioritize the design of scalable FL models that can handle an increasing number of clients without excessive computational and communication costs. This could involve the development of more efficient algorithms and the use of federated distillation techniques to reduce model size and complexity 137.
@@ 5.3.2 Negligence in Initialization
Most of the included studies began federated training from a random initialization, a method that, while effective in IID scenarios, can be less optimal for handling non-IID data. In healthcare, where data distribution often varies significantly across institutions due to differences in patient demographics, clinical practices, or data collection methods, random initialization can lead to slower convergence, increased communication costs, and potentially suboptimal local optima 138, 139.
A significant issue is the lack of standardization in model initialization approaches. Many studies either adopted random initialization without justification or entirely omitted the description of their initialization method. This inconsistency can result in significant variations in model performance and convergence rates, making it difficult to compare results across different studies and settings. Additionally, if the initial model is biased towards the data distribution of certain participants, it might not perform well across all clients, leading to fairness issues and suboptimal overall performance. Moreover, these challenges are further exacerbated by the heterogeneity of computational resources available across institutions. Some advanced initialization methods, such as those involving foundation models or pretraining on large-scale datasets, may be computationally expensive and thus infeasible for resource-constrained participants 138, 96, 43.
Personalization in model initialization is another underexplored area. Personalized initialization techniques, which tailor the starting point to the specific data distribution of each client, are critical for improving local model performance and accelerating convergence. However, research into these techniques, such as model-agnostic meta-learning and partial initialization for finding a good initialization, remains limited within FL for healthcare 140, 141.
@@
  • Transfer Learning for Initialization: Utilize pretraining or foundation models for initialization can provide a strong starting point for FL. This approach has been shown to not only speed up convergence but also mitigate the effects of both data and system heterogeneity 96, 43, 138, potentially closing the performance gap between FL and centralized learning 139. However, researchers should carefully consider the similarity between the source and target datasets to avoid negative transfer effects.
  • Personalized Initialization: Implement personalized initialization methods, such as model-agnostic meta-learning and partial initialization, can help customize the starting point of the model based on local data characteristics, further enhancing model performance on individual clients and improving overall system convergence 140, 141.
  • Standardization of Initialization Procedures: To ensure consistency and comparability across studies, there should be a push towards standardizing initialization methods in FL research. Providing clear descriptions and justifications for the chosen initialization strategy is essential for reproducibility and for advancing the field.
  • Consideration of Resource Constraints: When selecting initialization methods, researchers should account for the varying computational resources across participants. Techniques that balance initialization quality with computational feasibility, such as hierarchical model training or using lightweight pretrained models, are critical to ensuring broader applicability of FL in healthcare 96, 43, 138.
@@ 5.4.1 Heterogeneity Issues
Statistical heterogeneity arises due to the non-IID nature of healthcare data across various institutions, which is characterized by demographic differences, instrumentation biases, distinct data acquisition protocols, and human operations, etc 142. For instance, variations in CT scan quality across sites can lead to inconsistencies in the correlation between imaging data and corresponding site-specific EHR data. These inconsistencies severely degrade FL performance, with accuracy drops of up to 50% 4, necessitating additional communication rounds for convergence 143. Statistical heterogeneity can also result in clients overfitting to their local data, leading to poor generalization on data from other clients, making simple parameter averaging an ineffective aggregation strategy 144. Since local models are optimized towards different local optima, the aggregated global model may drift from the true global optima, causing a biased minimum and significantly slowing down convergence as illustrated in Figure 7. In healthcare, statistical heterogeneity can be broadly characterized by four forms, including:
@@
  • Quantity Skew. The number of training samples differs greatly across clients, leading to imbalanced data distributions. Models tend to optimizing for clients with more data, potentially neglecting those with less, further reducing the generalization ability 134.
@@
  • Label Skew. The distribution of labels varies across clients. For instance, in the context of COVID-19, hospitals in regions heavily impacted by the pandemic may have a higher proportion of positive cases, whereas other regions might have predominantly negative cases. This label imbalance can lead to biased models 145.
@@
  • Feature Skew. Different clients may have access to different features for the same sample cohort. For example, some institutions might only have access to EHR data, while others may have additional imaging modalities like X-rays or MRIs 140, 17. This is especially common in VFL scenarios.
@@
  • Quality Skew. This arises from varying data quality across clients, often due to issues like label noise, data acquisition noise, or processing discrepancies. Clients with high-quality, accurately labeled data contribute more effectively to model learning, while those with noisy or inaccurate labels can introduce errors, undermining the model’s generalization and convergence 146.
@@
Figure 8 provides an illustration of these skew forms. Numerous studies have been focusing on one of these skews in healthcare. 98 introduced a generative replay method by employing a VAE to synthesize medical images, enabling clients to train on a combined dataset of real and generated data, thus mitigating data quantity skew. 67 employed a frequency-based approach for medical data harmonization in FL, where images are processed in the frequency domain to retain local phase information and synchronize amplitudes across clients, thereby mitigating feature skew. 131 leveraged a Dirichlet distribution for modeling categorical probabilities in medical data, applying uncertainty calibration and diversity relaxation to enhance label annotation by focusing on samples with high uncertainty and low similarity, thus reducing label quality skew.
@@
Figure 8: Illustration of different skew forms in statistical heterogeneity, including quantity, feature, label, and quality skew. Medical image sources: Chest X-rays 147, CT scans 148, Ultrasound images 149, and Whole Slide Imaging (WSI) 150.
@@
Model heterogeneity occurs when different clients use varying model architectures due to differences in hardware capabilities, data types, or specific institutional requirements. For example, one hospital may use a complex deep learning model for image analysis, while another uses a simpler model due to computational constraints. This divergence complicates the process of aggregating model updates in FL, as different architectures may have different fitting capabilities, performance characteristics, and requirements for convergence 151. To address model heterogeneity, approaches like knowledge distillation have been employed, where a “student” model is used to transfer knowledge from various “teacher” models in different clients 151. However, these methods often require auxiliary public datasets for transferring knowledge, raising privacy concerns and increasing the computational burden on resource-constrained institutions 152, 151. An alternative approach is the use of a lightweight “messenger” model, as proposed by 99, which carries concentrated information from one client to another, reducing the need for a full model exchange. This method allows for efficient aggregation and distribution of knowledge without the overhead of auxiliary public datasets.
@@
System heterogeneity refers to the differences in computational capabilities, network architectures, and resource availability across clients. For example, some hospitals may have advanced computing infrastructure, while others may have limited hardware resources. This variability may affect the efficiency of FL, particularly in aggregating models trained on non-IID data 134, 136. Most studies have focused on addressing statistical or model heterogeneity, but system heterogeneity is equally important. Differences in hardware resources can lead to discrepancies in how models perform and converge across clients. Furthermore, the need for additional local computation and storage resources, as required by methods like knowledge distillation, can be a burden for institutions with limited resources 27, 152.
@@
  • Data Harmonization: Apply data harmonization techniques locally at each client site to minimize variability in data distributions and improve the consistency of FL models 67, 153.
@@
  • Feature Alignment and Data Imputation: Use methods like feature alignment and data imputation during training to ensure models learn from consistent data across clients. For example, FedHealth 77 demonstrated the ability to infer missing modalities in healthcare data through FTL.
  • Bias Checks and Corrections: Implement continuous monitoring for bias within and across clients throughout the FL process. Correct identified biases to prevent model performance degradation. For instance, 131 utilized uncertainty calibration and diversity relaxation to dynamically correct annotations for high-uncertainty, low-similarity samples.
  • Advanced Optimization Techniques: Enhance FL robustness against client drift and heterogeneity by employing advanced optimization techniques such as FedProx 154, SCAFFOLD 155, and MOON 156.
  • Disentangled Representation Learning (RDL): Integrate DRL into federated models to disentangle underlying invariant factors, making models more robust to data heterogeneity across different institutions 76, 157.
  • Fine-Tuning and Personalization: Improve model performance on domain-specific tasks by fine-tuning models locally with client-specific data and annotations. Extend techniques from meta-learning and multi-task learning to support personalized or device-specific modeling in FL 140.
@@
  • Adaptive FL Frameworks: Future research should focus on developing more adaptive FL frameworks capable of accommodating a wide range of model architectures and system configurations.

### v2

@@ 5.2.1 Unclear in Data Standardization and Harmonization
Healthcare data are often collected and stored in diverse and proprietary formats that do not always adhere to international standards and terminologies, complicating data linkage and reuse. For example, structured clinical data usually contains features that vary with differences in clinical practice across institutions 125, such as diabetes diagnosis, which can involve different glucose measurement methods with varying cut-off points, resulting in hidden heterogeneity that may be overlooked in subsequent statistical analyses. Additionally, language differences across institutions, especially in multilingual regions like the European Union, pose additional challenges in standardizing and harmonizing data. Medical terminology and clinical reports may be documented in different languages, complicating data interpretation and analysis across borders. A crucial step before implementing FL in healthcare is to ensure data standardization, harmonization, and interoperability across different cohorts, which are key to the success of FL (Figure 6).
Most simulation studies processed data centrally and generate artificially partitioned datasets without considering the distributed nature of various data silos. This oversight extends to the lack of discussion on how datasets at each client are curated for use in experiments. Despite this, almost all FL frameworks assume the input data is preformatted for model training or preprocessing pipelines. This assumption leads to significant frustration and delays, as the burden of data export and conversion typically falls on clinical data managers who may lack the necessary budget and training. Moreover, among the included studies, only two performed quality or integrity checks on the data. 126 excluded samples with impossible values (e.g., negative heart rates) and inconsistent feature values, while 127 used Principal Component Analysis to filter out noise. Few studies addressed structural or informative missingness, with only 29 and 40 considering imputation methods while also deleting features with high missingness rates. Poor quality imputation and handling of non-random missingness can bias model training. Additionally, no studies considered language differences in medical terminology and clinical reports across borders.
Figure 6: Illustration of Data Standardization and Harmonization. Multi-site T1-weighted MRI images from different scanners (GE, Philips, Siemens) exhibit variability in intensity distributions. The standardization process adjusts individual distributions to a common scale, while harmonization ensures alignment across datasets for improved consistency 128.
@@
  • Use Consistent Protocols: Collaborating clients must use consistent preprocessing protocols to standardize data to a Common Data Model (CDM) such as the Observational Medical Outcomes Partnership (OMOP) 129. Harmonizing healthcare data to a CDM like OMOP ensures it is interoperable with other clinical datasets, enabling effective merging and analysis across distributed sources and platforms 94.
@@
  • Address Language Differences: In multilingual regions like the European Union, language differences in medical terminology can hinder data interpretation. Translation and normalization techniques, along with Large Language Models (LLMs), can assist in automatic translation and ensuring consistent terminology, thereby improving data interoperability for FL applications.
@@ 5.2.2 Issues in Data Partition
Among all included studies, only a few leveraged natural splits to replicate data collection processes across different hospitals or institutions. For instance, 32 employed the CheXpert dataset, 130 worked with the chest X-ray dataset and 41 extracted metadata from Tissue Source sites in the TCGA dataset for their studies. These datasets naturally reflect the heterogeneity found in real-world clinical data across hospitals and institutions, making them more suitable for FL studies in healthcare 131, 59.
Simulation studies typically used heuristics to artificially create heterogeneous data partitions from a pooled dataset, assigning these partitions to simulated clients, as illustrated in Figure 5. Common synthetic partitioning methods for classification tasks include assigning samples from a limited number of classes to each client, using Dirichlet distribution sampling on class labels, and employing the Pachinko Allocation Method (PAM) when labels have a hierarchical structure 4. For regression tasks, Gaussian Mixture clustering based on t-SNE feature representations has been used to partition datasets among clients 132.
Nonetheless, synthetic partitioning methods may not accurately reflect the intricate heterogeneity found in real-world scenarios 91. Examples from digital histopathology illustrate the limitations of synthetic partitioning methods 133. In digital histopathology, tissue samples are extracted, stained, and digitized, leading to data heterogeneity due to factors such as patient demographics, staining techniques, physical slide storage methods, and digitization processes. Although advancements in staining normalization have reduced some heterogeneity, other sources remain challenging to replicate synthetically, and some may even be unknown 134. These underscore the necessity of conducting cohort experiments with natural splits to ensure that FL models are robust across varied clinical settings. This issue also extends to other areas, including radiology, dermatology, and retinal image analysis.
@@
Another significant issue is the lack of clear definitions and descriptions for train and test set partitions across clients in many studies. Among the studies included in this review, 84% did not explicitly define how these partitions are handled for each client, leading to potential ambiguity in evaluating model performance. This concern is particularly critical in the context of personalized FL, where each client’s test set should be unique to accurately reflect individual data distributions.
@@
  • Complement Simulation Studies with Real-World Data Evaluations: While simulation studies using artificially partitioned datasets can provide valuable insights, it is essential to validate these findings through evaluations on real-world, naturally partitioned datasets. This multi-stage evaluation process ensures that models are tested in both controlled environments and realistic deployment scenarios, improving their generalizability and robustness 135.
@@
  • Incorporate Multiple Types of Skew: Researchers should broaden the scope of their synthetic partitioning strategies to include not just quantity skew but also label skew and feature skew. This would create a more realistic representation of the heterogeneity found in real-world datasets, allowing FL models to be more robust and generalizable across diverse clinical settings 136.
  • Retaining Client-Specific Test Sets: Consider the non-IID nature of data in FL, we suggest that researchers retain a portion of data within each client as a dedicated test set rather than relying on a single, global test set for all clients. This approach provides a more accurate and reliable evaluation, reflecting the unique data distributions of each client, which is particularly important in personalized FL scenarios 135.
@@ 5.3.1 Inadequate Model Selection and Development
  • Standardization for Research and Benchmarking: Standardizing the process of model selection and development is essential for FL research, particularly in benchmarking and comparative studies. This standardization does not aim to limit innovation but to provide a consistent framework for evaluating models across diverse FL applications an settings. For instance, establishing shared protocols for selecting baseline models, defining performance metrics, and validating results can significantly enhance reproducibility and comparability. Such a framework encourages innovations that are both rigorous and generalizable, enabling the development of practical solutions tailored to the unique challenges of FL, such as data heterogeneity and resource constraints 54.
  • Exploring Simpler, More Interpretable Models: Researchers should not overlook simpler, traditional ML models, particularly in scenarios where interpretability is crucial. These models can be more practical and equally effective in certain healthcare applications, providing a balance between performance and interpretability 137, 21. Researchers should also consider developing simpler and lightweight models that can be deployed in resource-constrained environments without sacrificing performance.
  • Enhancing Personalization in FL: More work is needed to develop and refine personalized FL approaches that can adapt to the diverse data distributions encountered in healthcare. This could involve hybrid models that combine the strengths of both global and local models, as well as more sophisticated techniques for model adaptation 51.
  • Addressing Model Heterogeneity: Given the diverse requirements and constraints across different institutions, it is crucial to develop FL strategies that can effectively manage model heterogeneity. This includes exploring federated ensemble learning methods, which allow the aggregation of heterogeneous models and can lead to more robust and accurate predictions 138.
  • Improving Scalability and Efficiency: Future studies should prioritize the design of scalable FL models that can handle an increasing number of clients without excessive computational and communication costs. This could involve the development of more efficient algorithms and the use of federated distillation techniques to reduce model size and complexity 139.
@@ 5.3.2 Negligence in Initialization
Most of the included studies began federated training from a random initialization, a method that, while effective in IID scenarios, can be less optimal for handling non-IID data. In healthcare, where data distribution often varies significantly across institutions due to differences in patient demographics, clinical practices, or data collection methods, random initialization can lead to slower convergence, increased communication costs, and potentially suboptimal local optima 140, 141.
A significant issue is the lack of standardization in model initialization approaches. Many studies either adopted random initialization without justification or entirely omitted the description of their initialization method. This inconsistency can result in significant variations in model performance and convergence rates, making it difficult to compare results across different studies and settings. Additionally, if the initial model is biased towards the data distribution of certain participants, it might not perform well across all clients, leading to fairness issues and suboptimal overall performance. Moreover, these challenges are further exacerbated by the heterogeneity of computational resources available across institutions. Some advanced initialization methods, such as those involving foundation models or pretraining on large-scale datasets, may be computationally expensive and thus infeasible for resource-constrained participants 140, 97, 45.
Personalization in model initialization is another underexplored area. Personalized initialization techniques, which tailor the starting point to the specific data distribution of each client, are critical for improving local model performance and accelerating convergence. However, research into these techniques, such as model-agnostic meta-learning and partial initialization for finding a good initialization, remains limited within FL for healthcare 142, 143.
@@
  • Transfer Learning for Initialization: Utilize pretraining or foundation models for initialization can provide a strong starting point for FL. This approach has been shown to not only speed up convergence but also mitigate the effects of both data and system heterogeneity 97, 45, 140, potentially closing the performance gap between FL and centralized learning 141. However, researchers should carefully consider the similarity between the source and target datasets to avoid negative transfer effects.
  • Personalized Initialization: Implement personalized initialization methods, such as model-agnostic meta-learning and partial initialization, can help customize the starting point of the model based on local data characteristics, further enhancing model performance on individual clients and improving overall system convergence 142, 143.
  • Standardization of Initialization Procedures: Standardizing initialization in FL is essential for ensuring consistency, reproducibility, and comparability across studies. It provides a common baseline for benchmarking and reduces variability in outcomes 141. For example, Nguyen et al. 140 showed that consistent initialization improves convergence rates, especially under non-IID data, while pretraining-based methods help address system heterogeneity 97. Such standardization does not hinder innovation but establishes a foundation for exploring advanced techniques like meta-learning and hybrid initialization, ensuring their broader applicability in diverse healthcare settings.
  • Consideration of Resource Constraints: When selecting initialization methods, researchers should account for the varying computational resources across participants. Techniques that balance initialization quality with computational feasibility, such as hierarchical model training or using lightweight pretrained models, are critical to ensuring broader applicability of FL in healthcare 97, 45, 140.
@@ 5.4.1 Heterogeneity Issues
Statistical heterogeneity arises due to the non-IID nature of healthcare data across various institutions, which is characterized by demographic differences, instrumentation biases, distinct data acquisition protocols, and human operations, etc 144. For instance, variations in CT scan quality across sites can lead to inconsistencies in the correlation between imaging data and corresponding site-specific EHR data. These inconsistencies severely degrade FL performance, with accuracy drops of up to 50% 4, necessitating additional communication rounds for convergence 145. Statistical heterogeneity can also result in clients overfitting to their local data, leading to poor generalization on data from other clients, making simple parameter averaging an ineffective aggregation strategy 146. Since local models are optimized towards different local optima, the aggregated global model may drift from the true global optima, causing a biased minimum and significantly slowing down convergence as illustrated in Figure 7. In healthcare, statistical heterogeneity can be broadly characterized by four forms, including:
@@
  • Quantity Skew. The number of training samples differs greatly across clients, leading to imbalanced data distributions. Models tend to optimizing for clients with more data, potentially neglecting those with less, further reducing the generalization ability 136.
@@
  • Label Skew. The distribution of labels varies across clients. For instance, in the context of COVID-19, hospitals in regions heavily impacted by the pandemic may have a higher proportion of positive cases, whereas other regions might have predominantly negative cases. This label imbalance can lead to biased models 147.
@@
  • Feature Skew. Different clients may have access to different features for the same sample cohort. For example, some institutions might only have access to EHR data, while others may have additional imaging modalities like X-rays or MRIs 142, 17. This is especially common in VFL scenarios.
@@
  • Quality Skew. This arises from varying data quality across clients, often due to issues like label noise, data acquisition noise, or processing discrepancies. Clients with high-quality, accurately labeled data contribute more effectively to model learning, while those with noisy or inaccurate labels can introduce errors, undermining the model’s generalization and convergence 148.
@@
Figure 8 provides an illustration of these skew forms. Numerous studies have been focusing on one of these skews in healthcare. 99 introduced a generative replay method by employing a VAE to synthesize medical images, enabling clients to train on a combined dataset of real and generated data, thus mitigating data quantity skew. 68 employed a frequency-based approach for medical data harmonization in FL, where images are processed in the frequency domain to retain local phase information and synchronize amplitudes across clients, thereby mitigating feature skew. 133 leveraged a Dirichlet distribution for modeling categorical probabilities in medical data, applying uncertainty calibration and diversity relaxation to enhance label annotation by focusing on samples with high uncertainty and low similarity, thus reducing label quality skew.
@@
Figure 8: Illustration of different skew forms in statistical heterogeneity, including quantity, feature, label, and quality skew. Medical image sources: Chest X-rays 149, CT scans 150, Ultrasound images 151, and Whole Slide Imaging (WSI) 152.
@@
Model heterogeneity occurs when different clients use varying model architectures due to differences in hardware capabilities, data types, or specific institutional requirements. For example, one hospital may use a complex deep learning model for image analysis, while another uses a simpler model due to computational constraints. This divergence complicates the process of aggregating model updates in FL, as different architectures may have different fitting capabilities, performance characteristics, and requirements for convergence 153. To address model heterogeneity, approaches like knowledge distillation have been employed, where a “student” model is used to transfer knowledge from various “teacher” models in different clients 153. However, these methods often require auxiliary public datasets for transferring knowledge, raising privacy concerns and increasing the computational burden on resource-constrained institutions 154, 153. An alternative approach is the use of a lightweight “messenger” model, as proposed by 100, which carries concentrated information from one client to another, reducing the need for a full model exchange. This method allows for efficient aggregation and distribution of knowledge without the overhead of auxiliary public datasets.
@@
System heterogeneity refers to the differences in computational capabilities, network architectures, and resource availability across clients. For example, some hospitals may have advanced computing infrastructure, while others may have limited hardware resources. This variability may affect the efficiency of FL, particularly in aggregating models trained on non-IID data 136, 138. Most studies have focused on addressing statistical or model heterogeneity, but system heterogeneity is equally important. Differences in hardware resources can lead to discrepancies in how models perform and converge across clients. Furthermore, the need for additional local computation and storage resources, as required by methods like knowledge distillation, can be a burden for institutions with limited resources 27, 154.
@@
  • Data Harmonization: Apply data harmonization techniques locally at each client site to minimize variability in data distributions and improve the consistency of FL models 68, 155.
@@
  • Feature Alignment and Data Imputation: Use methods like feature alignment and data imputation during training to ensure models learn from consistent data across clients. For example, FedHealth 78 demonstrated the ability to infer missing modalities in healthcare data through FTL.
  • Bias Checks and Corrections: Implement continuous monitoring for bias within and across clients throughout the FL process. Correct identified biases to prevent model performance degradation. For instance, 133 utilized uncertainty calibration and diversity relaxation to dynamically correct annotations for high-uncertainty, low-similarity samples.
  • Advanced Optimization Techniques: Enhance FL robustness against client drift and heterogeneity by employing advanced optimization techniques such as FedProx 156, SCAFFOLD 157, and MOON 158.
  • Disentangled Representation Learning (RDL): Integrate DRL into federated models to disentangle underlying invariant factors, making models more robust to data heterogeneity across different institutions 77, 159.
  • Fine-Tuning and Personalization: Improve model performance on domain-specific tasks by fine-tuning models locally with client-specific data and annotations. Extend techniques from meta-learning and multi-task learning to support personalized or device-specific modeling in FL 142.
@@
  • Adaptive FL Frameworks: Future research should focus on developing more adaptive FL frameworks capable of accommodating a wide range of model architectures and system configurations. Examples of adaptive frameworks include FedAdapt 160, which adjusted client participation and resource allocation based on system heterogeneity, and AutoFL 161, which used AutoML techniques to dynamically configure FL processes. Adaptive frameworks can also be realized through dynamic client selection and hierarchical aggregation (e.g., Clustered FL 162) to balance data quality, system resources, and communication efficiency.